# An accurate mean-field equation for voter model dynamics on scale-free networks

### v3

@@ Abstract
Understanding the emergent macroscopic behavior of dynamical systems on networks is a crucial but challenging task. One of the simplest and most effective methods to construct a reduced macroscopic model is given by mean-field theory. The resulting approximations perform well on dense and homogeneous networks but poorly on scale-free networks, which, however, are more realistic in many applications. In this paper, we introduce a modified version of the mean-field approximation for voter model dynamics on scale-free networks. The two main deviations from classical theory are that we use degree-weighted shares as coarse variables and that we introduce a correlation factor that can be interpreted as slowing down dynamics induced by interactions. We observe that the correlation factor is only a property of the network and not of the state or of parameters of the process. This approach achieves a significantly smaller approximation error than standard methods without increasing dimensionality.
@@ I Introduction
The voter model [Clifford1973, Holley1975] and its many extensions (see, e.g., [Castellano2009, Porter2016, Redner2019] for overviews) are among the most popular models for studying spreading or infection processes on networks. In this paper, we focus on a general variant, the continous-time noisy voter model (CNVM) [Luecke2023, Luecke2024a, Carro2016]. Given a simple graph $G$ on $N$ nodes, each node $i\in\{0,\dots,N-1\}=:[N]$ has a state $x_{i}\in\{0,\dots,M-1\}=:[M]$. As the voter model is often applied to opinion dynamics, we refer to $x_{i}$ as the opinion of node $i$. Each opinion $x_{i}$ evolves over time according to a continuous-time Markov chain: if $x_{i}$ is in state $a\in[M]$, it transitions to a different state $b\in[M]$ at the rate
@@
where $\mathbf{1}_{\{x_{i}=a\}}\in\{0,1\}$ denotes the indicator function of the event $x_{i}=a$. Using mean-field theory [Barrat2008], it has been shown that, in the large-population limit $N\to\infty$, the trajectory $\bm{c}(t)$ of time-dependent shares converges to the solution of a mean-field equation (MFE), provided that the underlying network is sufficiently homogeneous and dense [Luecke2023, Keliger2024, Gleeson2012, Ayi2021]. For scale-free networks—e.g., generated by the Barabási–Albert (BA) model [Albert2002] with preferential attachment parameter mm—however, this classical mean-field approximation performs poorly [Fernley2022, Gleeson2013, PastorSatorras2001]. The main contribution of this paper is the derivation and validation of a modified mean-field equation (mMFE) that can accurately capture the macroscopic behavior of the CNVM on scale-free networks.
It is well known [Luecke2023] that for the CNVM the MFE is a simple system of ordinary differential equations
@@
The significance of the degree-weighted shares is already known in some contexts [Bianconi2002, Suchecki2005], and they were specifically determined to be excellent collective variables for the CNVM on BA networks using a data-driven method developed by the authors [Luecke2024].
In this paper, we derive the above mMFE (5), motivating why these modifications are necessary for scale-free networks, and show how to obtain the correlation factor $\alpha$. We analyze the quality and robustness of the mMFE depending on the model parameters $r$ and $\tilde{r}$, using a theoretical error analysis as well as numerical studies. We show that the mMFE approximates the degree-weighted shares exceptionally well when the imitation rates $r_{a,b}$ are similarly large and there is moderate noise $\tilde{r}$. Moreover, in this setting the mMFE is robust in the sense that $\alpha$ does not depend on the specific model parameters and initial state—hence, the correlation factor appears to be a universal attribute of the network, more precisely of the network’s statistical properties. In other parameter regimes, the approximation quality degrades slightly. Due to its popularity, the numerical examples presented in the main text utilize the BA model to generate scale-free networks, but almost identical results can be obtained for other examples like the configuration model [Bollobas1980], as shown in Appendix B. Specifically for BA networks, another interesting result is that the correlation factor $\alpha$ increases towards $1$ with increasing preferential attachment parameter $m$ of the BA network. Finally, we show that the mMFE outperforms standard methods of similar complexity like the standard homogeneous MFE (3) and pair-approximation (PA) [Porter2016, Gleeson2013] significantly. While higher-dimensional methods like the heterogeneous MFE and PA [Porter2016, Pugliese2009], approximate master equation [gleeson2011high], quenched MFE and PA [Mata2013], or dynamic cavity method [Neri2009] are able to resolve more detailed information about the system than the mMFE, they also require substantially more variables. The mMFE is, on the other hand, able to provide a parsimonious macroscopic description of the collective dynamics using only a minimal set of variables: the degree-weighted shares $\bar{\bm{c}}$.
Accompanying code for reproducing the numerical experiments presented in this paper is provided at [lucke_2025_17122500].
@@ II.1 Relation to classical mean-field theory
Moreover, in the classical theory it is typically argued that nodes become interchangeable in the mean-field limit, so that the expected state of every node is simply given by the average state, e.g., $\mathbb{E}[x_{i}]=c$ for all $i$ in the binary-state case ($M=2$). We observe numerically that, for BA networks, this interchangeability is also valid on average; however, although the expected state of each node is related to the degree-weighted average state $\bar{c}$ instead, as described in assumption (14). As mentioned earlier, this finding was inspired by the authors’ work [Luecke2024], which computationally revealed that the weighted shares are excellent collective variables for the CNVM on BA networks. This means that their future evolution depends only on the current value of $\bar{c}$, and on no further information from the full state $\bm{x}=(x_{i})_{i\in[N]}$.
@@ II.2 Error analysis
(Note that $\frac{1}{2}\lVert\bar{\bm{c}}(t)-\bar{\bm{c}}_{\text{{\tiny mMFE}}}(t)\rVert_{1}\in[0,1]$ is the total variation distance between discrete probability distributions [Rachev1991].) Then, choosing the factor
@@ II.3 Parameter regimes
    Regime D: Dissimilar rates and (partially) no noise. Example: $r_{0,1}=1$, $r_{1,0}=0$, $\tilde{r}_{0,1}=0$, $\tilde{r}_{1,0}=0.3$ (i.e., SIS model [Kiss2017]).
@@ IV Comparison to classical approximations
In this section we compare the mMFE, which is a (M1)(M-1)-dimensional approximation, to other standard low-dimensional moment-closure approximations: the MFE, which also has dimension (M1)(M-1), and the pair-approximation (PA), which has a dimension of (M1)+M(M1)(M-1)+M(M-1). The PA results from closing the moment equations at second order and therefore includes the number of edges between different opinions as additional variables compared to the MFE. A detailed derivation of the PA for the CNVM is omitted here for brevity but the general idea as well as the resulting equations are presented in appendix C. As Fig. 6 illustrates, the classical approximations (MFE and PA) fail to capture the system behavior on BA networks, whereas the mMFE incurs only a very small error. This is not surprising, since the classical approximations assume a homogeneous network structure. The well-known heterogeneous variants of the MFE and PA, or the approximate master equation, are better suited for scale-free networks [gleeson2011high, Porter2016, Gleeson2013, Pugliese2009]. However, as these resolve the opinion shares per degree, they involve substantially more variables than the mMFE. For instance, on a BA network with $N=10^{4}$ nodes and preferential attachment parameter $m=4$, there are on average about 100 different node degrees, implying that the heterogeneous MFE has 100 times more variables than our mMFE. The quenched MFE [Mata2013] even resolves the expected state per node such that the required number of variables is $N$ times larger than that of the mMFE. Hence, to compare reduced order models with the same degree of efficiency, we only report comparison between the mMFE and the homogeneous MFE and PA.
@@ V Conclusion
The inspiration for the mMFE (5) arose from findings in [Luecke2024], where a computational approach was proposed to identify forecastable variables of dynamical processes on networks. For such processes, reduced modeling is typically guided by theoretical insights—our approach therefore constitutes an exception. Along these lines, it would be highly desirable to gain a theoretical understanding of the observed superior behavior of the mMFE. Furthermore, this work also highlights the importance of data-driven approaches, since many of the insights presented here were supported by computational analysis. These methods could be utilized in future work to derive low-dimensional approximations for other systems that are not yet well understood analytically.
@@ Appendix B Scale-free networks generated by the configuration model
In this section we verify that the mMFE (5) is also valid for scale-free networks generated by the configuration model [Bollobas1980], which can be used to generate a random graph with an arbitrary degree sequence $(d_{i})_{i\in[N]}$. The method works as follows: After assigning $d_{i}$ half-edges to each node $i$, two half-edges are repeatedly picked from all remaining half-edges, joined together to form an edge, and removed from the set of available half-edges, until all half-edges have been used up. This potentially yields a multigraph since self-loops and multiple edges between two nodes are possible. In such a case we simply restart and repeat the whole procedure until finding a simple graph. To produce a scale-free network, we first generate a degree sequence $(d_{i})_{i\in[N]}$ such that the associated degree distribution follows a power law with exponent $\gamma=3$, and then apply the configuration model. To facilitate a fair comparison to a BA model with preferential attachment parameter $m$, the minimum degree in the degree sequence should also be chosen as $m$.
Asymptotically, the BA model and the configuration model are relatively similar. For example, for both models the average clustering coefficient falls to zero rapidly at a rate of approximately $1/N$, and the diameter grows moderately at a rate of approximately $\log N$ [Luecke2024a]. Furthermore, the assortativity, i.e., the degree correlation between neighbors, is asymptotically zero for both models. However, for the finite network sizes of $N\approx 10^{4}$ considered here, the BA model and the configuration model exhibit significant differences. The most noteworthy aspect is that for the BA model the assortativity is still clearly negative, meaning that it is more likely for small degree nodes to be connected to large degree nodes, while for the configuration model the assortativity is already close to zero. Moreover, the BA model shows a considerably larger clustering coefficient than the configuration model.
@@ Appendix C Pair-approximation
In this section we briefly outline the derivation of the classical pair-approximation in the case of $M=2$ opinions. A more detailed derivation can readily be found in the literature, see for instance [Peralta2018, Luecke2024a]. Instead of closing the first-order moment equation using an independence assumption $\mathbb{E}[(1-x_{i})x_{j}]\approx\mathbb{E}[1-x_{i}]\mathbb{E}[x_{j}]$, the pair approximation assumes that all neighbor pairs are statistically equivalent and introduces the average pair probability $\mathbb{E}[(1-x_{i})x_{j}]=:s$ as an additional variable. Based on (II), this yields

### v4

# Accurate mean-field equation for voter model dynamics on scale-free networks

@@ Abstract
Understanding the emergent macroscopic behavior of dynamical systems on networks is a crucial but challenging task. One of the simplest and most effective methods to construct a reduced macroscopic model is given by mean-field theory. The resulting approximations perform well on dense and homogeneous networks but poorly on scale-free networks, which, however, are more realistic in many applications. In this paper, we introduce a modified version of the mean-field approximation for voter model dynamics on scale-free networks. The two main deviations from classical theory are that we use degree-weighted shares as coarse variables and that we introduce a correlation factor that can be interpreted as slowing down dynamics induced by interactions. We observe that, for moderate noise and comparable interaction rates, the correlation factor is only a property of the network and not of the state or of parameters of the process. This approach achieves a significantly smaller approximation error than standard methods without increasing dimensionality.
@@ I Introduction
The voter model [8, 13] and its many extensions (see, e.g., [7, 24, 27] for overviews) are among the most popular models for studying spreading or infection processes on networks. In this paper, we focus on a general variant, the continuous-time noisy voter model (CNVM) [16, 18, 6]. Given a simple graph $G$ on $N$ nodes, each node $i\in\{0,\dots,N-1\}=:[N]$ has a state $x_{i}\in\{0,\dots,M-1\}=:[M]$. As the voter model is often applied to opinion dynamics, we refer to $x_{i}$ as the opinion of node $i$. Each opinion $x_{i}$ evolves over time according to a continuous-time Markov chain: if $x_{i}$ is in state $a\in[M]$, it transitions to a different state $b\in[M]$ at the rate
@@
where $\mathbf{1}_{\{x_{i}=a\}}\in\{0,1\}$ denotes the indicator function of the event $x_{i}=a$. Using mean-field theory [3], it has been shown that, in the large-population limit $N\to\infty$, the trajectory $\bm{c}(t)$ of time-dependent shares converges to the solution of a mean-field equation (MFE), provided that the underlying network is sufficiently homogeneous and dense [16, 14, 10, 2]. For scale-free networks—e.g., generated by the Barabási–Albert (BA) model [1] with preferential attachment parameter mm—however, this classical mean-field approximation performs poorly [9, 11, 22]. The main contribution of this paper is the derivation and validation of a modified mean-field equation (mMFE) that can accurately capture the macroscopic behavior of the CNVM on scale-free networks.
It is well known [16] that for the CNVM the MFE is a simple system of ordinary differential equations
@@
The significance of the degree-weighted shares is already known in some contexts [4, 28], and they were specifically determined to be excellent collective variables for the CNVM on BA networks using a data-driven method developed by the authors [17].
In this paper, we derive the above mMFE (5), motivating why these modifications are necessary for scale-free networks, and show how to obtain the correlation factor $\alpha$. We analyze the quality and robustness of the mMFE depending on the model parameters $r$ and $\tilde{r}$, using a theoretical error analysis as well as numerical studies. We show that the mMFE approximates the degree-weighted shares exceptionally well when the imitation rates $r_{a,b}$ are similarly large and there is moderate noise $\tilde{r}$. Moreover, in this setting the mMFE is robust in the sense that $\alpha$ does not depend on the specific model parameters and initial state—hence, the correlation factor appears to be a universal attribute of the network, more precisely of the network’s statistical properties. In other parameter regimes, the approximation quality degrades slightly. Due to its popularity, the numerical examples presented in the main text utilize the BA model to generate scale-free networks, but almost identical results can be obtained for other examples like the configuration model [5], as shown in Appendix B. Specifically for BA networks, another interesting result is that the correlation factor $\alpha$ increases towards $1$ with increasing preferential attachment parameter $m$ of the BA network. Finally, we show that the mMFE outperforms standard methods of similar complexity like the standard homogeneous MFE (3) and pair-approximation (PA) [24, 11] significantly. While higher-dimensional methods like the heterogeneous MFE and PA [24, 25], approximate master equation [12], quenched MFE and PA [20], or dynamic cavity method [21] are able to resolve more detailed information about the system than the mMFE, they also require substantially more variables. The mMFE is, on the other hand, able to provide a parsimonious macroscopic description of the collective dynamics using only a minimal set of variables: the degree-weighted shares $\bar{\bm{c}}$.
Accompanying code for reproducing the numerical experiments presented in this paper is provided at [19].
@@ II.1 Relation to classical mean-field theory
Moreover, in the classical theory it is typically argued that nodes become interchangeable in the mean-field limit, so that the expected state of every node is simply given by the average state, e.g., $\mathbb{E}[x_{i}]=c$ for all $i$ in the binary-state case ($M=2$). We observe numerically that, for BA networks, this interchangeability is also valid on average; however, the expected state of each node is related to the degree-weighted average state $\bar{c}$ instead, as described in assumption (14). As mentioned earlier, this finding was inspired by the authors’ work [17], which computationally revealed that the weighted shares are excellent collective variables for the CNVM on BA networks. This means that their future evolution depends only on the current value of $\bar{c}$, and on no further information from the full state $\bm{x}=(x_{i})_{i\in[N]}$.
@@ II.2 Error analysis
(Note that $\frac{1}{2}\lVert\bar{\bm{c}}(t)-\bar{\bm{c}}_{\text{{\tiny mMFE}}}(t)\rVert_{1}\in[0,1]$ is the total variation distance between discrete probability distributions [26].) Then, choosing the factor
@@ II.3 Parameter regimes
    Regime D: Dissimilar rates and (partially) no noise. Example: $r_{0,1}=1$, $r_{1,0}=0$, $\tilde{r}_{0,1}=0$, $\tilde{r}_{1,0}=0.3$ (i.e., SIS model [15]).
@@ IV Comparison to classical approximations
In this section we compare the mMFE, which is a (M1)(M-1)-dimensional approximation, to other standard low-dimensional moment-closure approximations: the MFE, which also has dimension (M1)(M-1), and the pair-approximation (PA), which has a dimension of (M1)+M(M1)(M-1)+M(M-1). The PA results from closing the moment equations at second order and therefore includes the number of edges between different opinions as additional variables compared to the MFE. A detailed derivation of the PA for the CNVM is omitted here for brevity but the general idea as well as the resulting equations are presented in appendix C. As Fig. 6 illustrates, the classical approximations (MFE and PA) fail to capture the system behavior on BA networks, whereas the mMFE incurs only a very small error. This is not surprising, since the classical approximations assume a homogeneous network structure. The well-known heterogeneous variants of the MFE and PA, or the approximate master equation, are better suited for scale-free networks [12, 24, 11, 25]. However, as these resolve the opinion shares per degree, they involve substantially more variables than the mMFE. For instance, on a BA network with $N=10^{4}$ nodes and preferential attachment parameter $m=4$, there are on average about 100 different node degrees, implying that the heterogeneous MFE has 100 times more variables than our mMFE. The quenched MFE [20] even resolves the expected state per node such that the required number of variables is $N$ times larger than that of the mMFE. Hence, to compare reduced order models with the same degree of efficiency, we only report comparison between the mMFE and the homogeneous MFE and PA.
@@ V Conclusion
The inspiration for the mMFE (5) arose from findings in [17], where a computational approach was proposed to identify forecastable variables of dynamical processes on networks. For such processes, reduced modeling is typically guided by theoretical insights—our approach therefore constitutes an exception. Along these lines, it would be highly desirable to gain a theoretical understanding of the observed superior behavior of the mMFE. Furthermore, this work also highlights the importance of data-driven approaches, since many of the insights presented here were supported by computational analysis. These methods could be utilized in future work to derive low-dimensional approximations for other systems that are not yet well understood analytically.
@@ Appendix B Scale-free networks generated by the configuration model
In this section we verify that the mMFE (5) is also valid for scale-free networks generated by the configuration model [5], which can be used to generate a random graph with an arbitrary degree sequence $(d_{i})_{i\in[N]}$. The method works as follows: After assigning $d_{i}$ half-edges to each node $i$, two half-edges are repeatedly picked from all remaining half-edges, joined together to form an edge, and removed from the set of available half-edges, until all half-edges have been used up. This potentially yields a multigraph since self-loops and multiple edges between two nodes are possible. In such a case we simply restart and repeat the whole procedure until finding a simple graph. To produce a scale-free network, we first generate a degree sequence $(d_{i})_{i\in[N]}$ such that the associated degree distribution follows a power law with exponent $\gamma=3$, and then apply the configuration model. To facilitate a fair comparison to a BA model with preferential attachment parameter $m$, the minimum degree in the degree sequence should also be chosen as $m$.
Asymptotically, the BA model and the configuration model are relatively similar. For example, for both models the average clustering coefficient falls to zero rapidly at a rate of approximately $1/N$, and the diameter grows moderately at a rate of approximately $\log N$ [18]. Furthermore, the assortativity, i.e., the degree correlation between neighbors, is asymptotically zero for both models. However, for the finite network sizes of $N\approx 10^{4}$ considered here, the BA model and the configuration model exhibit significant differences. The most noteworthy aspect is that for the BA model the assortativity is still clearly negative, meaning that it is more likely for small degree nodes to be connected to large degree nodes, while for the configuration model the assortativity is already close to zero. Moreover, the BA model shows a considerably larger clustering coefficient than the configuration model.
@@ Appendix C Pair-approximation
In this section we briefly outline the derivation of the classical pair-approximation in the case of $M=2$ opinions. A more detailed derivation can readily be found in the literature, see for instance [23, 18]. Instead of closing the first-order moment equation using an independence assumption $\mathbb{E}[(1-x_{i})x_{j}]\approx\mathbb{E}[1-x_{i}]\mathbb{E}[x_{j}]$, the pair approximation assumes that all neighbor pairs are statistically equivalent and introduces the average pair probability $\mathbb{E}[(1-x_{i})x_{j}]=:s$ as an additional variable. Based on (II), this yields
@@ Appendix D Supplementary figures
In this section, two supplementary figures are provided. Figure S3 shows the dependence of the mMFE error on the difference of imitation rates, and Fig. S4 shows the mMFE error for varying numbers of nodes.